OpenStudy (anonymous):

Solve the following equations for x. Identify any extraneous solutions. Show all work. -4 sqrt(x+9)=20 I think what I did was wrong. I got x=567

OpenStudy (anonymous):

\[-4\sqrt{x+9}=20\]

OpenStudy (anonymous):

-4*sqrt(x+9) = 20 Divide both sides by -4 and you get: sqrt(x + 9) = -5 Then square both sides and you get X + 9 = 25 Minus 9 from both sides and you get: x = 16 Then test 16. -4*sqrt(16+9) = 20 -4*5 = -20 I believe x = 16 is an extraneous root.

OpenStudy (anonymous):

-4 * 5 = -20 which is not equal to 20, thats why I think x =16 is an extraneous root.

OpenStudy (anonymous):

thanks